OpenStudy (anonymous):

The sum of three consecutive integers is 33. Find the integers

OpenStudy (anonymous):

Show work

OpenStudy (nurali):

x+(x+1)+(x+2)=33 3x+3=33 Subtract both sides by 3 3x+3-3=33-3 3x=30 Divide both side by 3 3x/3=30/3 x=10

OpenStudy (nurali):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (anonymous):

Wait isn't there more to the answer?

OpenStudy (anonymous):

it's 10,11 and 12 @sabby75488 :)

OpenStudy (anonymous):

3 consecutive integers are 10,11 and 12

OpenStudy (nurali):

x=10 put value in this equation x+(x+1)+(x+2)=33 10+11+12=33 33=33

OpenStudy (anonymous):

whose sum is 33. :)

OpenStudy (anonymous):

Okay can we show work idk how you got 10,11, and 12?

OpenStudy (anonymous):

@Nurali had show u the working. :)

OpenStudy (nurali):

x=10 answer value x+1=10+1=11 x+2=10+2=12

OpenStudy (anonymous):

\[x+(x+1)+(x+2)=33\]\[x+x+1+x+2=33\]\[3x+3=33\]\[3x=33-3\]\[3x=30\]\[x=10\]substitute x=10 in the equation \[x+(x+1)+(x+2)=33\]\[(10)+(10+1)+(10+2)=33\]\[10+11+12=33\]Therefore the 3 consecutive integers are 10,11 and 12 whose sum is 33. @sabby75488